### v1

# A Fourier-RKHS approach for characterizing equivalent Gaussian distributions for stationary processes on homogeneous spaces

Abstract
Pairs of equivalent Gaussian distributions for centered stationary processes on homogeneous spaces can be characterized in terms of their spectral measures. The purpose of this note is to consider the latter characterization from the perspective of a reproducing kernel Hilbert space (RKHS) approach.
Keywords: Equivalent Gaussian distributions for stationary processes; reproducing kernel Hilbert spaces; spherical functions
2020 MSC: 60G10; 60G15; 60G30; 46E22; 43A90
1 Introduction
1.1 Notational conventions
We refer to $\mathbb{N}$ as the set of strictly positive integers. In the absence of ambiguity, we write $(a_{n})=(a_{n})_{n\in\mathbb{N}}$ for a sequence indexed over $\mathbb{N}$. Given a topological space $(E,\tau)$, the Borel $\sigma$-field over $E$ is denoted by $\mathfrak{B}(E)$. The space of continuous functions $f$ on $E$ is written as $C(E)$. Also, $f\in C_{c}(E)$ if and only if $f\in C(E)$ and $f$ has compact support. For $x,y\in\mathbb{R}^{d}$, $\langle x,y\rangle=x^{\mathrm{t}}y$ identifies the dot product on $\mathbb{R}^{d}$. Given a measure space $(E,\mathcal{A},\mu)$, the space of measurable functions $f\colon E\to\mathbb{C}$ that are square integrable on $E$ w.r.t. $\mu$ is denoted with $L^{2}(E,\mathcal{A},\mu)$. The canonical norm of $\langle f,g\rangle_{\mu}=\int_{E}f(x)\overline{g(x)}\mu(dx)$ on $L^{2}(E,\mathcal{A},\mu)$ is denoted with $\lVert\cdot\rVert_{\mu}$. If clear from the context, we make use of the short notations $L^{2}(\mu)$ or $L^{2}(E)$. The space of absolutely integrable functions (on $E$ w.r.t. $\mu$) is written as $L^{1}(E)$. Given two measures $\mu_{1}$ and $\mu_{1}$ on $\mathcal{A}$ and $L^{2}(\mu_{1})\supset L\subset L^{2}(\mu_{2})$, we use the notation $\lVert\cdot\rVert_{\mu_{1}}\asymp\lVert\cdot\rVert_{\mu_{2}}$ on $L$ to indicate that the norms $\lVert\cdot\rVert_{\mu_{1}}$ and $\lVert\cdot\rVert_{\mu_{2}}$ are equivalent on $L$. That is, there exist constants $\alpha_{1},\alpha_{2}>0$, s.t. for any $\varphi\in L$, $0<\alpha_{1}\lVert\varphi\rVert_{\mu_{2}}\leq\lVert\varphi\rVert_{\mu_{1}}\leq\alpha_{2}\lVert\varphi\rVert_{\mu_{2}}<\infty$. The measures $\mu_{1}$ and $\mu_{2}$ are termed equivalent on $\mathcal{A}$ if they are mutually absolutely continuous on $\mathcal{A}$, i.e., $\mu_{1}(A)=0$ implies $\mu_{2}(A)=0$, $A\in\mathcal{A}$, and vice versa. If $\mu_{1}$ and $\mu_{2}$ are equivalent on $\mathcal{A}$ we write $\mu_{1}\equiv\mu_{2}$ on $\mathcal{A}$. On the other hand, $\mu_{1}$ and $\mu_{2}$ are referred to as orthogonal on $\mathcal{A}$, written as $\mu_{1}\perp\mu_{2}$ on $\mathcal{A}$, if there exists a separating set $A\in\mathcal{A}$ for which $\mu_{1}(A)=0$ and $\mu_{2}(\Omega\setminus A)=0$.
1.2 Symmetric positive-definite kernels and their RKHS
Let $T$ be a set. A function $R\colon T\times T\to\mathbb{R}$ is referred to as a symmetric nonnegative-definite kernel if $R(s,t)=R(t,s)$ for all $s,t\in T$, and if for any $n\in\mathbb{N}$ and $t_{1},\dotsc,t_{n}\in T$,
$\sum_{i=1}^{n}\sum_{j=1}^{n}a_{i}a_{j}R(t_{i},t_{j})\geq 0,\quad a_{1},\dotsc,a_{n}\in\mathbb{R}.$ (1)
If equality in (1) holds only for $a_{1}=\cdots=a_{n}=0$, $R$ is said to be strictly positive-definite. Given a symmetric nonnegative-definite kernel $R$, we write $H_{T}(R)$ for the Hilbert space of real-valued functions on $T$ which satisfies
$R(\cdot,t)\in H_{T}(R),\quad t\in T,$
and
$f(t)=\langle f,R(\cdot,t)\rangle_{R},\quad t\in T,\quad f\in H_{T}(R),$
where $\langle\cdot,\cdot\rangle_{R}$ denotes the inner product on $H_{T}(R)$. This identifies $H_{T}(R)$ as the unique RKHS with reproducing kernel $R$. The existence and uniqueness statement regarding $H_{T}(R)$ is known as the Moore–Aronszajn theorem (cf. [Aronszajn, RKHSBook]). For $n\in\mathbb{N}$ and $T_{n}=\{t_{1},\dotsc,t_{n}\}\subset T$, we write $R(n)$ for the $n\times n$ matrix with entries $R(t_{i},t_{j})$. If $R$ is strictly positive-definite, one can identify $H_{T_{n}}(R(n))=\mathbb{R}^{n}$ with inner product
$\langle v,w\rangle_{R(n)}=v^{\mathrm{t}}R(n)^{-1}w,\quad v,w\in\mathbb{R}^{n}.$
Example 1.1.
Let $(E,\mathcal{A},\mu)$ be a measure space with $\sigma$-finite measure $\mu$. Assume that $\gamma(t,\cdot)\in L^{2}(E,\mathcal{A},\mu)$, $t\in T$. Denote with $L_{T}(\gamma)$ the closed subspace of $L^{2}(E,\mathcal{A},\mu)$ spanned by $\{\gamma(t,\cdot)\colon t\in T\}$. Let $R$ be a symmetric nonnegative-definite kernel s.t.
$R(s,t)=\int_{E}\!\gamma(s,u)\overline{\gamma(t,u)}\mu(du),\quad s,t\in T.$
Then, $H_{T}(R)$ consists of real-valued functions $f(t)=\int_{E}\!\xi_{f}(u)\gamma(t,u)\mu(du)$, $\xi_{f}\in L_{T}(\gamma)$, with inner product $\langle f,g\rangle_{R}=\langle\xi_{f},\xi_{g}\rangle_{\mu}$.
Remark 1.1.
In the following, it is assumed that any RKHS $H_{T}(R)$ is separable. Recall that if $T$ is a topological space and $R$ is continuous on $T\times T$, the separability of $H_{T}(R)$ is equivalent to the separability of $T$. In particular, if nothing else is mentioned, any topological space $T$ is assumed to be separable.
1.3 Gaussian processes and their equivalent distributions
Let $X=(X_{t})_{t\in T}$ be a real-valued and centered Gaussian process, defined on a probability space $(\Omega,\mathcal{F},P_{1})$, with covariance function $R_{1}$. The $\sigma$-field generated by $X$ is denoted by $\sigma_{T}(X)$. Let $P_{2}$ be another probability measure on $\sigma_{T}(X)$ s.t. the process $X$ under $P_{2}$ is real-valued, centered, and Gaussian, with covariance function $R_{2}$. In particular, for $\ell=1,2$, $P_{\ell}$ is referred to as a centered Gaussian measure on $\sigma_{T}(X)$ with covariance function $R_{\ell}$. The linear span of $\{X_{t}\colon t\in T\}$ is written as $H_{0}(X)$ and we denote by $H_{\ell}(X)$ its closure in $L^{2}(P_{\ell})$. That is, $H_{\ell}(X)$ is the Gaussian space associated with the process $X$ under $P_{\ell}$.
Example 1.2.
Let $W$ be the Wiener measure on the Borel $\sigma$-field over $C([0,1])$. Introduce the transformation $g_{\sigma_{\ell}}(x)=\sigma_{\ell}x$, $x\in C([0,1])$, where $\sigma_{\ell}$ is a real number, strictly positive. Define the measure $P_{\ell}$ on $\sigma_{[0,1]}(X)$ by
$P_{\ell}(A)=W(g_{\sigma_{\ell}}^{-1}(B)),\quad A=X^{-1}(B),\ B\in\mathfrak{B}(C([0,1])).$
It follows that $X$ under $P_{\ell}$ is a Brownian motion with covariance function $R_{\ell}(s,t)=\sigma^{2}\operatorname{min}\{s,t\}$.
A valuable consideration is that $P_{1}$ and $P_{2}$ are either equivalent or orthogonal. This dichotomy has been verified by various authors (cf. [ParzenTimeSeries], p. 478). Feldman [Feldman1958] provides a characterization in terms of the Gaussian spaces $H_{1}(X)$ and $H_{2}(X)$. In particular, he shows that for $P_{1}$ and $P_{2}$ to be equivalent on $\sigma_{T}(X)$, it is necessary and sufficient that there exists a linear homeomorphism $T$ from $H_{1}(X)$ onto $H_{2}(X)$ s.t. $T^{*}T-I$ is Hilbert-Schmidt. He refers to $T$ as an equivalence operator from $H_{1}(X)$ onto $H_{2}(X)$. Later, Rozanov [Rozanov] provides an alternative proof using the entropy of the measure $P_{1}$ w.r.t. $P_{2}$. An approach which was pioneered earlier by Hájek [HajekJDiv, HajekOriginal]. Let $T_{n}=\{t_{1},\dotsc,t_{n}\}$ be a finite collection of coordinates from $T$ and denote by $\sigma(Y_{n})$ the $\sigma$-field generated by the Gaussian vector $Y_{n}=(X_{t_{1}},\dotsc,X_{t_{n}})$. Also, let $P_{\ell}^{n}$ be the restriction of $P_{\ell}$ to $\sigma(Y_{n})$ and write $\mathcal{K}$ for the class of all finite subsets of $T$. Then, Hájek [HajekOriginal] shows that
$\text{$P_{1}\equiv P_{2}$ on $\sigma_{T}(X)$}\quad\text{ if and only if}\quad\text{$\sup_{T_{n}\in\mathcal{K}}J(n)<\infty$},$ (2)
where,
$J(n)=\begin{cases}E_{2}\Big[\operatorname{log}\frac{dP_{2}^{n}}{dP_{1}^{n}}\Big]-E_{1}\Big[\operatorname{log}\frac{dP_{2}^{n}}{dP_{1}^{n}}\Big],&\text{if $P_{1}^{n}\equiv P_{2}^{n}\text{ on }\sigma(Y_{n})$,}\\ \infty,&\text{otherwise.}\end{cases}$ (3)
Notice that if $R_{1}$ and $R_{2}$ are strictly positive-definite, then $\nicefrac{{dP_{2}^{n}}}{{dP_{1}^{n}}}=\nicefrac{{p_{2}^{n}(Y_{n})}}{{p_{1}^{n}(Y_{n})}}$ with
$p_{\ell}^{n}(y)=\frac{\operatorname{exp}\big(-\frac{1}{2}\langle y,y\rangle_{R_{\ell}(n)}\big)}{\sqrt{(2\pi)^{n}\det R_{\ell}(n)}},\quad y\in\mathbb{R}^{n}.$ (4)
The number (3) is known as the J-divergence (or just divergence) between the finite-dimensional distributions $P_{1}^{n}$ and $P_{2}^{n}$ of $X$ (cf. [Jeffreys] p. 158 and [KullbackLeibler]).
Example 1.3.
A particularly simple case is $R_{1}$ and $R_{2}$ strictly positive-definite and s.t. $R_{1}=\alpha^{2}R_{2}$, $\alpha\neq 1$. In this case,
$J(n)=\frac{1}{2}\bigg(\alpha-\frac{1}{\alpha}\bigg)^{2}n,$ (5)
which can be deduced by evaluating
$2J(n)=\operatorname{tr}(A_{n}^{12})+\operatorname{tr}(A_{n}^{21})-2n,\quad A_{n}^{ij}=R_{i}(n)[R_{j}(n)]^{-1},\quad i,j=1,2.$ (6)
If we reconsider Example 1.2 and take $X$ as a Brownian motion with covariance function $R_{\ell}(s,t)=\sigma_{\ell}^{2}\min\{s,t\}$ under $P_{\ell}$ ($\sigma_{1}\neq\sigma_{2}$), it follows from (5) that $P_{1}\perp P_{2}$ on $\sigma_{[0,1]}(X)$. This result can be traced back to the work of Cameron and Martin [CameronMartin] — compare also with Example 1 in Hájek’s paper [HajekOriginal].
1.4 RKHS characterization of equivalent Gaussian distributions
Given two Hilbert spaces $\mathcal{H}_{1}$ and $\mathcal{H}_{2}$ of functions defined on $T$, write $u\mapsto\widetilde{u}$ for the mapping defined by,
$\widetilde{u}(s,t)=\sum_{k=1}^{n}f_{k}(s)g_{k}(t),\quad u=\sum_{k=1}^{n}f_{k}\otimes g_{k},$ (7)
where for any $k=1,\dotsc,n$, $f_{k}\otimes g_{k}\in\mathcal{H}_{1}\otimes\mathcal{H}_{2}$ is an elementary tensor. In view of the RKHS choices $H_{T}(R_{1})$ and $H_{T}(R_{2})$, we let $H_{T\times T}(R_{1}\otimes R_{2})$ denote the RKHS on $T\times T$ with kernel $R_{1}\otimes R_{2}\big((s_{1},s_{2}),(t_{1},t_{2})\big)=R_{1}(s_{1},t_{1})R_{2}(s_{2},t_{2})$. Then, it is known that the map $u\mapsto\widetilde{u}$ extends to an isometry between the Hilbert spaces $H_{T}(R_{1})\otimes H_{T}(R_{2})$ and $H_{T\times T}(R_{1}\otimes R_{2})$. This is summarized by writing
$\widetilde{H_{T}(R_{1})\otimes H_{T}(R_{2})}=H_{T\times T}(R_{1}\otimes R_{2}).$ (8)
A proof of the latter equality is given by Aronszajn [Aronszajn] (cf. Section 8). We also refer to Section 5.5 in [RKHSBook]. Early work on the RKHS characterization of equivalent Gaussian distributions was done by Parzen [ParzenRKHS], and Kallianpur and Oodaira [KallianpurOodaira1] (cf. Oodaira [Oodaira]). For instance, for $T$ countably infinite, Parzen [ParzenRKHS] relies on Hájek’s characterization of equivalence (2) and proclaims the necessity of $R_{2}-R_{1}\in H_{T\times T}(R_{1}\otimes R_{2})$ for $P_{1}$ and $P_{2}$ to be equivalent on $\sigma_{T}(X)$. Furthermore, an expression of the Radon–Nikodym derivative of $P_{2}$ w.r.t. $P_{1}$ is derived (cf. Capon [Capon]). Later, a general result is obtained by Neveu [Neveu]. For simplicity, we write $H_{T}(R_{\ell})^{2\otimes}=H_{T}(R_{\ell})\otimes H_{T}(R_{\ell})$, $R_{\ell}^{2\otimes}=R_{\ell}\otimes R_{\ell}$ and denote by $H_{T}(R_{\ell})^{2\odot}=H_{T}(R_{\ell})\odot H_{T}(R_{\ell})$ the closed subspace of $H_{T}(R_{\ell})^{2\otimes}$ composed of symmetric tensors. Accordingly, let $R_{\ell}^{2\odot}=R_{\ell}\odot R_{\ell}$ be defined by
$R_{\ell}^{2\odot}[(s_{1},s_{2}),(t_{1},t_{2})]=\frac{1}{2}\big(R_{\ell}(s_{1},t_{1})R_{\ell}(s_{2},t_{2})+R_{\ell}(s_{1},t_{2})R_{\ell}(s_{2},t_{1})\big).$
In particular, using the isometric correspondence between $H_{T}(R_{\ell})^{2\otimes}$ and $H_{T\times T}(R_{\ell}^{2\otimes})$, it follows that $H_{T}(R_{\ell})^{2\odot}$ and $H_{T\times T}(R_{\ell}^{2\odot})$ are isometric. Let $\langle\cdot,\cdot\rangle_{2\odot}$ denote the inner product on $H_{T}(R_{1})^{2\odot}$ and write $\mathscr{U}\colon H_{T}(R_{1})\to H_{T}(R_{1})^{*}$ for the Hilbert-Schmidt operator $\mathscr{U}(f)(g)=\langle f\odot g,U\rangle_{2\odot}$ associated with an element $U\in H_{T}(R_{\ell})^{2\odot}$. Then, using the inherent structure of the Gaussian space $H_{1}(X)$, Neveu [Neveu] (cf. Proposition 8.6) shows that $P_{1}$ and $P_{2}$ are equivalent on $\sigma_{T}(X)$ if and only if there exists $u\in H_{T\times T}(R_{1}^{2\odot})$ s.t.
$R_{2}(s,t)-R_{1}(s,t)=\langle R_{1}^{2\odot}\big[(s,t),\cdot\big],u\rangle_{R_{1}^{2\odot}}\ \text{($\Leftrightarrow$ $R_{2}-R_{1}\in H_{T\times T}(R_{1}^{2\odot})$)},$ (9)
and the eigenvalues of the Hilbert-Schmidt operator $\mathscr{U}$, associated with the corresponding element $H_{T}(R_{1})^{2\odot}\ni U$ of $u$, are strictly larger than 1-1. Notice that (9) is equivalent to $R_{2}-R_{1}\in H_{T\times T}(R_{1}^{2\otimes})$. We also point out that for $P_{1}$ and $P_{2}$ to be equivalent on $\sigma_{T}(X)$ it is necessary that $H_{T}(R_{1})=H_{T}(R_{2})$. This can be derived from the correspondence $H_{T}(R_{1})=H_{T}(R_{2})$ if and only if $\lVert\cdot\rVert_{P_{1}}\asymp\lVert\cdot\rVert_{P_{2}}$ on $H_{0}(X)$, which is a consequence of Aronszajn’s differences of kernels theorem (cf. [Aronszajn], Corollary IV3 on p. 383). In particular, we can substitute the assumption on the operator $\mathscr{U}$ and obtain the following RKHS characterization of equivalent Gaussian distributions (cf. Chatterji and Mandrekar [Chatterji_Mandrekar]): $P_{1}\equiv P_{2}$ on $\sigma_{T}(X)$ if and only if
  1. (a)
    $H_{T}(R_{1})=H_{T}(R_{2})$;
  2. (b)
    $R_{2}-R_{1}\in H_{T\times T}(R_{1}^{2\otimes})$ and $0\equiv m\in H_{T}(R_{1})$.
Example 1.4 (Stationary processes on real coordinate spaces).
Let $T=\mathbb{R}^{d}$ and assume that
$R_{\ell}(s+h,t+h)=R_{\ell}(s,t),\quad s,t,h\in\mathbb{R}^{d}.$
That is, $X$ is stationary under $P_{\ell}$. Define $k_{\ell}(t)=R_{\ell}(t,0)$, $t\in\mathbb{R}^{d}$. We observe that $R_{\ell}(s,t)=k_{\ell}(s-t)$. In particular, the assignment $(s,t)\mapsto k_{\ell}(s-t)$ is a symmetric nonnegative-definite kernel. Assume that $k_{\ell}$ is continuous at zero. By Bochner’s theorem (for real coordinate spaces) [Bochner],
$R_{\ell}(s,t)=\int_{\mathbb{R}^{d}}\!\operatorname{e}^{{i\mkern 1.0mu}\langle s,\lambda\rangle}\overline{\operatorname{e}^{{i\mkern 1.0mu}\langle t,\lambda\rangle}}F_{\ell}(d\lambda),\quad s,t\in\mathbb{R}^{d},$
for some finite measure $F_{\ell}$, uniquely defined on $\mathfrak{B}(\mathbb{R}^{d})$. Notice that we are in the framework of Example 1.1. Actually, if $R_{\ell}$ is strictly positive-definite and $F_{\ell}(d\lambda)=\eta_{\ell}(\lambda)d\lambda$, with spectral density $\eta_{\ell}$, it follows that $H_{\mathbb{R}^{d}}(R_{\ell})$ consists of continuous functions $f\in L^{1}(\mathbb{R}^{d})\cap L^{2}(\mathbb{R}^{d})$ s.t. $\nicefrac{{\hat{f}}}{{\sqrt{\hat{k}_{\ell}}}}\in L^{2}(\mathbb{R}^{d})$ with inner product,
$\langle f,g\rangle_{R_{\ell}}=(2\pi)^{\nicefrac{{-d}}{{2}}}\int_{\mathbb{R}^{d}}\!\frac{\hat{f}(\lambda)\overline{\hat{g}(\lambda)}}{\hat{k}_{\ell}(\lambda)}d\lambda.$
In the latter, $h\mapsto\hat{h}$ denotes the Fourier transform of $h$ (cf. Wendland [Wendland_Book], Theorem 10.12). In particular, $H_{\mathbb{R}^{d}}(R_{\ell})\subset L^{2}(\mathbb{R}^{d})$. Upon the isometric correspondence given in (8), it follows that $H_{\mathbb{R}^{d}\times\mathbb{R}^{d}}(R_{\ell}^{2\otimes})\subset L^{2}(\mathbb{R}^{d}\times\mathbb{R}^{d})$. Then, since the mapping
$t\mapsto\int_{\mathbb{R}^{d}}\!\delta(s,t)^{2}ds,\quad\delta(s,t)=k_{2}(s-t)-k_{1}(s-t),$
is constant, we conclude that $R_{2}-R_{1}$ does not belong to $H_{\mathbb{R}^{d}\times\mathbb{R}^{d}}(R_{1}^{2\otimes})$, unless $R_{1}$ and $R_{2}$ are equal. This shows that for $T=\mathbb{R}^{d}$, the Gaussian measures $P_{1}$ and $P_{2}$ are orthogonal on $\sigma_{\mathbb{R}^{d}}(X)$ as soon as $R_{1}\neq R_{2}$. A recent treatment — for $T$ not necessarily equal to $\mathbb{R}^{d}$ but sufficiently dense — is given in [hediger]. If $T$ is bounded, the situation is different. As an example, Striebel [Striebel] (cf. [Capon]) gives an expression of the Radon–Nikodym derivative of $P_{2}$ w.r.t. $P_{1}$ for the case where $T=[0,b]$ and
$R_{1}(s,t)=\sigma^{2}_{1}\operatorname{e}^{-\beta_{1}\lvert s-t\rvert},\quad\quad R_{2}(s,t)=\sigma^{2}_{2}\operatorname{e}^{-\beta_{2}\lvert s-t\rvert},\quad 2\sigma^{2}_{1}\beta_{1}=2\sigma^{2}_{2}\beta_{2}.$
As for a collection of results concerning the characterization of equivalent Gaussian distributions for real stationary processes, an overview is given in the books by Yadrenko [Yadrenko], Ibragimov and Rozanov [GaussianRandomProcesses], and Gikhman and Skorokhod [StochasticProcesses].
Example 1.5 (Isotropic processes on the sphere).
Let $T=\mathbb{S}^{d-1}$, $d\geq 3,$ be the unit sphere in $\mathbb{R}^{d}$. Denote with $\Delta\colon\mathbb{S}^{d-1}\times\mathbb{S}^{d-1}\to[0,\pi]$ the great-circle (geodesic) distance on $\mathbb{S}^{d-1}$, i.e., $\Delta(s,t)=\operatorname{arccos}(\langle s,t\rangle)$. Assume that
$R_{\ell}(s,t)=\psi_{\ell}(\Delta(s,t)),\quad s,t\in\mathbb{S}^{d-1},$
where $\psi_{\ell}\colon[0,\pi]\to\mathbb{R}$ is continuous and s.t. $\psi_{\ell}(0)>0$. Given $k\in\mathbb{N}\cup\{0\}$, write $S_{k}^{j}$, $j=1,\dotsc,h(k)$, for the spherical harmonics of degree $k$ (cf. Chapter XI (Section 11.3) in [Erdelyi] or also Chapter IV (Section 2) in [SteinWeiss]). Then, in the sense of Schoenberg [Schoenberg], the following series representation of $R_{\ell}$ is valid,
$R_{\ell}(s,t)=\sum_{k=0}^{\infty}\sum_{j=1}^{h(k)}S_{k}^{j}(s)S_{k}^{j}(t)a_{\ell}(k).$ (10)
In the latter, the coefficients $a_{\ell}(k)$ are strictly positive for infinitely many $k$ (cf. p. 72 in [Yadrenko] or also Theorem 1 in [GneitingMultiquadric]). An explicit description of $H_{\mathbb{S}^{d-1}}(R_{\ell})$ is given in [RKHSSphere]. In particular, any member $f$ of $H_{\mathbb{S}^{d-1}}(R_{\ell})$ is linked with a square summable sequence $(c_{k,j}(f))$, $j=1,\dotsc,h(k)$, s.t.
$f(t)=\sum_{k=0}^{\infty}\sum_{j=1}^{h(k)}S_{k}^{j}(t)c_{k,j}(f)\sqrt{a_{\ell}(k)},\quad\langle f,g\rangle_{R_{\ell}}=\sum_{k=0}^{\infty}\sum_{j=1}^{h(k)}c_{k,j}(f)c_{k,j}(g).$ (11)
Assume w.l.o.g. that $R_{2}-R_{1}\in H_{\mathbb{S}^{d-1}\times\mathbb{S}^{d-1}}(R_{1}^{2\otimes})$ is derived from an elementary tensor, i.e., there exist $f,g\in H_{\mathbb{S}^{d-1}}(R_{1})$ s.t. $(R_{2}-R_{1})(s,t)=f(s)g(t)$. Upon the description given in (11), it follows that
$(R_{2}-R_{1})(s,t)=\sum_{k=0}^{\infty}\sum_{j=1}^{h(k)}\sum_{p=0}^{\infty}\sum_{q=1}^{h(p)}S_{k}^{j}(s)S_{p}^{q}(t)c_{k}^{j}(f)c_{p}^{q}(g)\sqrt{a_{2}(k)a_{2}(p)}.$
Comparing coefficients with (10), we observe that
$c_{k}^{j}(f)c_{p}^{q}(g)=\begin{cases}\frac{a_{2}(k)-a_{1}(k)}{a_{2}(k)},&k=p,\\ 0,&\text{otherwise,}\end{cases}$
and
$\lVert R_{2}-R_{1}\rVert_{R_{1}^{2\otimes}}=\sum_{k=0}^{\infty}h(k)\bigg(1-\frac{a_{1}(k)}{a_{2}(k)}\bigg)^{2}.$ (12)
Therefore, $P_{1}\equiv P_{2}$ on $\sigma_{\mathbb{S}^{d-1}}(X)$ if and only if the latter sum is finite. Upon the results from Feldman and Hájek, one can arrive at the same conclusion (cf. Theorem 1 in [GaussianSphereEquivOrtho]).
1.5 Stationary processes on locally compact abelian groups
1.5.1 Positive-definite functions
Let $G$ be a locally compact group. In general, we write $g\cdot\tilde{g}=g\tilde{g}$ for the group operation and $e=gg^{-1}$ identifies the identity element on $G$. A continuous function $\varphi\colon G\to\mathbb{C}$ is said to be positive-definite if for any $n\in\mathbb{N}$ and $g_{1},\dotsc,g_{n}\in G$,
$\sum_{i=1}^{n}\sum_{j=1}^{n}z_{i}\overline{z}_{j}\varphi(g_{i}g_{j}^{-1})\geq 0,\quad z_{1},\dotsc,z_{n}\in\mathbb{C}.$
If $G$ is abelian, i.e., $G$ is a locally compact abelian (LCA), we denote the group operation by ++ and the identity element is written as 0.
1.5.2 Positive-definite functions on LCA groups
Let $G$ be a LCA group. Recall that $G^{*}$ consists of continuous characters of $G$, i.e., $\chi\in G^{*}$ if and only if $\chi\colon G\to\mathbb{C}$ is continuous, $\lvert\chi(g)\rvert=1$, $g\in G$, and
$\chi(g+\tilde{g})=\chi(g)\chi(\tilde{g}),\quad g,\tilde{g}\in G.$
In particular, $\chi(0)=1$ and $\chi(-g)=\overline{\chi(g)}$, $g\in G$. The following is known as Bochner’s theorem [Rudin]. A continuous function $\phi\colon G\to\mathbb{C}$ is positive-definite if and only if,
$\phi(g)=\int_{G^{*}}\chi(g)\mu(d\chi),\quad g\in G,$ (13)
for some finite measure $\mu$, uniquely defined on $\mathfrak{B}(G^{*})$.
Example 1.6.
If $G=\mathbb{R}^{d}$, the dual group $G^{*}$ is isomorphic to $\mathbb{R}^{d}$ with isomorphism $\lambda\mapsto\operatorname{exp}({i\mkern 1.0mu}\langle\cdot,\lambda\rangle)$. In particular, any continuous character $\chi$ of $\mathbb{R}^{d}$ is given by $\chi(\cdot)=\chi_{\lambda}(\cdot)=\operatorname{exp}({i\mkern 1.0mu}\langle\cdot,\lambda\rangle)$ for some $\lambda\in\mathbb{R}^{d}$. Upon the identification $T\colon G^{*}\to\mathbb{R}^{d}$, $T(\chi_{\lambda})=\lambda$, it follows from (13) that any continuous positive-definite function $\phi\colon\mathbb{R}^{d}\to\mathbb{C}$ admits a representation $\phi(x)=\int_{\mathbb{R}^{d}}\!\operatorname{exp}({i\mkern 1.0mu}\langle x,\lambda\rangle)F(d\lambda)$ where $F$ is the pushforward of $\mu$ by $T$.
1.5.3 Equivalent Gaussian distributions on LCA groups
Let $T=G$. Assume that $X$ is stationary under $P_{\ell}$. That is, for any triple $s,t,g\in G$, $R_{\ell}(s+g,t+g)=R_{\ell}(s,t)$. It follows that the function $k_{\ell}\colon G\to\mathbb{C}$, defined by $k_{\ell}(t)=R_{\ell}(t,0)$, $t\in G$, is positive-definite. Suppose that $k_{\ell}$ is continuous at 0. We deduce from (13) that there exists a finite measure $\mu_{\ell}$ on $\mathfrak{B}(G^{*})$ s.t.
$R_{\ell}(s,t)=\int_{G^{*}}\chi(s)\overline{\chi(t)}\mu_{\ell}(d\chi),\quad s,t\in G.$
Then, the following characterization of equivalent Gaussian distributions on LCA groups is due to Grenander [Grenander] (cf. [Chatterji_Mandrekar]). For $P_{1}$ and $P_{2}$ to be equivalent on $\sigma_{T}(X)$ it is necessary and sufficient that the nonatomic parts of $\mu_{1}$ and $\mu_{2}$ agree and $\mu_{1}$ and $\mu_{2}$ share the same set of atoms $\{a_{n}\colon n\in\mathbb{N}\}$ s.t.
$\sum_{n\in\mathbb{N}}\bigg(1-\frac{\mu_{1}(\{a_{n}\})}{\mu_{2}(\{a_{n}\})}\bigg)^{2}<\infty.$
1.6 Bi-invariant functions of positive type
1.6.1 Spherical characterization
[…]
$\varphi(g)=\int_{\mathscr{L}}\!\zeta(g)\mu(d\zeta).$ (14)
The latter representation is due to Godement [Godement1958] (cf. the earlier work of Gelfand [Gelfand]).
1.6.2 Translation to invariant kernels on homogeneous spaces
Let $T=G/K$, the set of left cosets of $K$ in $G$. Assume that $R_{\ell}$ is $G$-invariant, i.e., $X$ is stationary under $P_{\ell}$ — for any $g\in G$ and $s,t\in T$, $R_{\ell}(gs,gt)=R_{\ell}(s,t)$. It follows that $R_{\ell}(s,t)=R_{\ell}(g_{s}^{-1}g_{t}K,K)$, $s,t\in T$, $s=g_{s}K$, $t=g_{t}K$. Therefore, if we define $\varphi_{\ell}(g)=R_{\ell}(gK,K)$, $\varphi_{\ell}$ is bi-invariant and positive-definite. We refer to $\varphi_{\ell}$ as the $K$-invariant version of $R_{\ell}$. If $\varphi_{\ell}$ happens to be continuous at the identity, we deduce from (14) that
$R_{\ell}(s,t)=\varphi_{\ell}(g_{s}^{-1}g_{t})=\int_{\mathscr{L}}\!\zeta(g_{s}^{-1}g_{t})\mu_{\ell}(d\zeta),\quad s=g_{s}K,\ t=g_{t}K.$ (15)
Remark 1.2.
Given the choices for $G$ and $K$, besides the requirements for $R_{\ell}$ (resp. $\varphi_{\ell}$), the only assumption underlying (15) is that $(G,K)$ is a Gelfand pair. Specifically, if $G$ is a Lie group, any (Riemannian) symmetric pair $(G,K)$ is a Gelfand pair. This result is due to Gelfand [Gelfand] (cf. [Helgason], p. 408). In the context of invariant stochastic processes, having in mind the appearance of (15), further reading is given in the book by Malyarenko [Malyarenko] (cf. Yaglom [YaglomInvariant]).
Remark 1.3.
We recall that if $T^{\prime}$ is a $G$ set with transitive group action $g\cdot t=gt$ and $K=K_{p}$ is the stabilizer subgroup of $G$ w.r.t. some $p\in T^{\prime}$, then $T^{\prime}$ and $T=G/K$ are isomorphic. For $R_{\ell}$ defined on $T^{\prime}\times T^{\prime}$, invariant w.r.t. the group action $G$, (15) becomes $R_{\ell}(s^{\prime},t^{\prime})=\int_{\mathscr{L}}\!\zeta(g_{s^{\prime}}^{-1}g_{t^{\prime}})\mu_{\ell}(d\zeta)$, with $s^{\prime}=g_{s^{\prime}}p$ and $t^{\prime}=g_{t^{\prime}}p$. In this setting, it can be assumed w.l.o.g. that $T^{\prime}=T$.
Example 1.7 (Rotation invariance).
If $G=SO(d)$, $d\in\mathbb{N}$, the covariance function given in Example 1.5, $R_{\ell}(s,t)=\psi_{\ell}(\Delta(s,t))$, is $G$ invariant — its representation based on spherical harmonics is an instance of (15) with Gelfand pair (SO(d),SO(d1))(SO(d),SO(d-1)).
Example 1.8.
Let $M$ be a Riemannian (globally) symmetric space. Denote by $G$ the identity component of the isometry group of $M$ and let $K=K_{p}$ be the stabilizer subgroup of $G$ w.r.t. an arbitrary point $p$ of $M$. Then, $(G,K)$ is a symmetric pair.
1.6.3 Equivalent Gaussian distributions on homogeneous spaces
Given $\zeta\in\mathscr{L}$, it is always possible to find a corresponding (spherical) irreducible unitary representation $\pi_{\zeta}$ on a Hilbert space $H_{\zeta}$ with cyclic vector $u_{\zeta}$. That is, for any $g\in G$, $\zeta(g)=\langle\pi_{\zeta}(g)u_{\zeta},u_{\zeta}\rangle_{\zeta}$, where $\langle\cdot,\cdot\rangle_{\zeta}$ denotes the inner product on $H_{\zeta}$ (see for instance Theorem 3.4, Chapter IV, §3, in [Helgason]). If $H_{\zeta}$ is finite dimensional, $d(\zeta)$ denotes the dimension of $H_{\zeta}$, otherwise $d(\zeta)=\infty$. The following is due to Chow [chow1, chow2] [^1] Chow’s result is stated more generally. It allows for less symmetric $T$, such as $G$-spaces possessing a dense orbit. We have formulated the statement in the context of Gelfand pairs $(G,K)$.. Under the assumption of (15), $P_{1}\equiv P_{2}$ on $\sigma_{T}(X)$ if and only if
  1. (i)
    the nonatomic parts of $\mu_{1}$ and $\mu_{2}$ agree;
  2. (ii)
    $\mu_{1}$ and $\mu_{2}$ share the same set of atoms $\{a_{n}\colon n\in\mathbb{N}\}$ s.t.
    $\sum_{n\in\mathbb{N}}d(a_{n})\bigg(1-\frac{\mu_{1}(\{a_{n}\})}{\mu_{2}(\{a_{n}\})}\bigg)^{2}<\infty.$
The proof of Chow is based on Feldman’s [Feldman1958] characterization of equivalent Gaussian distributions. In the following, we will give an RKHS interpretation of Chow’s result.
2 Equivalent Gaussian distributions on homogeneous spaces: A Fourier-RKHS approach
Let $T=G/K$, where $G$ and $K$ are as in the previous section. In particular, $G$ is unimodular locally compact and $K$ a compact subgroup of $G$ s.t. $(G,K)$ is a Gelfand pair. Assume that $R_{\ell}$ is $G$-invariant with $K$-invariant version $\varphi_{\ell}$ that is continuous at the identity. Let $\varphi_{\ell}\in L^{1}(G)\cap L^{2}(G)$ — or equivalently, $R_{\ell}(\cdot,t)\in L^{1}(T)\cap L^{2}(T)$ for any $t\in T$. Given $f\in L^{1}(T)$, we write $\hat{f}(\zeta)$ for the Fourier transform of $f$ at $\zeta\in\mathscr{L}$ (cf. [Wolf], p. 196). Let $\mathcal{H}(\ell)$ be composed of continuous functions $f\in L^{1}(T)\cap L^{2}(T)$ which are of the form
$f(t)=\int_{\mathscr{L}}\!\big\langle\hat{f}(\zeta),\widehat{R_{\ell}(\cdot,t)}(\zeta)\big\rangle_{\zeta}\mu_{\scriptscriptstyle P}(d\zeta),$ (16)
where $\mu_{\scriptscriptstyle P}$ is the Plancherel measure for $(G,K)$ on $\mathscr{L}$. Since $R_{\ell}$ is real-valued, we accept the notational convenience $\operatorname{Re}f=f$. Then, if we equip $\mathcal{H}(\ell)$ with inner product
$\langle f_{1},f_{2}\rangle_{\mathcal{H}(\ell)}=\int_{\mathscr{L}}\!\big\langle\hat{f}_{1}(\zeta),\hat{f}_{2}(\zeta)\big\rangle_{\zeta}\mu_{\scriptscriptstyle P}(d\zeta),$ (17)
it follows from the Plancherel theorem (cf. Corollary 9.6.6 in [Wolf]) that $\mathcal{H}(\ell)$ is a Hilbert space. In particular, for any $t\in T$, $R_{\ell}(\cdot,t)\in\mathcal{H}(\ell)$ and $f(t)=\langle f,R_{\ell}(\cdot,t)\rangle_{\mathcal{H}(\ell)}$, $f\in\mathcal{H}(\ell)$. This is summarized in the following lemma.
Lemma.
For $\varphi_{\ell}\in L^{1}(G)\cap L^{2}(G)$, $H_{T}(R_{\ell})$ consists of functions $f\in L^{1}(T)\cap L^{2}(T)$ which are of the form (16) with inner product given by (17). Additionally, $\lVert f\rVert_{R_{\ell}}=\lVert f\rVert_{L^{2}(T)}$.
The previous lemma permits the following observations.
Non-compact case
For $\varphi_{\ell}\in L^{1}(G)\cap L^{2}(G)$ and $G$ not compact, $P_{1}\equiv P_{2}$ on $\sigma_{T}(X)$ if and only if $\mu_{1}=\mu_{2}$. To see it, notice that for any $v,w\in G$, the difference $R_{2}-R_{1}](vK,wK)$ is given by $\varphi_{2}-\varphi_{1}](v^{-1}w)$. We know that $P_{1}\equiv P_{2}$ on $\sigma_{T}(X)$ is equivalent to $R_{2}-R_{1}\in H_{T\times T}(R_{1}^{2\otimes})$. By the lemma, the latter is equivalent to $R_{2}-R_{1}\in L^{2}(T\times T)$, which is true if and only if $\varphi_{1}=\varphi_{2}$. Or equivalently, $\mu_{1}=\mu_{2}$ (by uniqueness of the measures $\mu_{1}$ and $\mu_{2}$).
Compact case
We recall that if $G$ is compact, $d(\zeta)<\infty$, $\zeta\in\mathscr{L}$, and the Plancherel measure $\mu_{\scriptscriptstyle P}$ is purely atomic (cf. Chapter 5 in [Wolf]). Also, by Mercer’s theorem (see for instance [RKHSBook]), $\varphi_{\ell}\in L^{2}(G)$. In particular, we are in the setting of the lemma. By construction of the Plancherel measure, since $\varphi_{\ell}$ is bi-invariant, positive-definite and a member of $C(G)\cap L^{1}(G)$, we have that $\mu_{\ell}(d\zeta)=\hat{\varphi}_{\ell}(\zeta)\mu_{\scriptscriptstyle P}(d\zeta)$. In the latter, $\hat{\varphi}_{\ell}(\zeta)$ is the spherical transform of $\varphi_{\ell}$ at $\zeta$ according to Theorem 9.4.1 in [Wolf]. Therefore, $\mu_{\ell}$ is purely atomic. Also, by uniqueness of the Plancherel measure on $\mathscr{L}$, $\mu_{1}$ and $\mu_{2}$ have the same set of atoms $\{a_{n}\colon n\in\mathbb{N}\}$. In particular, up to $\mu_{\scriptscriptstyle P}$ measure zero,
$\hat{\varphi}_{1}(\zeta)=\frac{\mu_{1}(d\zeta)}{\mu_{2}(d\zeta)}\hat{\varphi}_{2}(\zeta)\quad\Leftrightarrow\quad\widehat{\varphi_{2}-\varphi_{1}}(\zeta)=\bigg(1-\frac{\mu_{1}(d\zeta)}{\mu_{2}(d\zeta)}\bigg)\hat{\varphi}_{2}(\zeta).$ (18)
By the lemma, $P_{1}$ and $P_{2}$ are equivalent on $\sigma_{T}(X)$ if and only if
$\big[R_{2}-R_{1}\big](s,t)=\int_{\mathscr{L}^{2}}\!\big\langle\widehat{R_{2}-R_{1}}(\zeta_{1},\zeta_{2}),\widehat{R_{1}^{2\otimes}\big[\cdot,(s,t)\big]}(\zeta_{1},\zeta_{2})\big\rangle_{\zeta_{1}\otimes\zeta_{2}}\mu_{\scriptscriptstyle P}\otimes\mu_{\scriptscriptstyle P}[d(\zeta_{1},\zeta_{2})],$
where
$\int_{\mathscr{L}^{2}}\!\lVert\widehat{R_{2}-R_{1}}(\zeta_{1},\zeta_{2})\rVert^{2}_{\zeta_{1}\otimes\zeta_{2}}\mu_{\scriptscriptstyle P}\otimes\mu_{\scriptscriptstyle P}[d(\zeta_{1},\zeta_{2})]<\infty.$ (19)
Then, since $R_{\ell}$ is $G$-invariant, it follows from the uniqueness of the Fourier transform that
$\widehat{R_{2}-R_{1}}(\zeta_{1},\zeta_{2})=\widehat{\varphi_{2}-\varphi_{1}}(\zeta_{1})\mathbbm{1}_{\{\zeta_{1}\}}(\zeta_{2}).$
By Fubini, we deduce from (18) and (19) that
$\int_{\mathscr{L}}\!\Big\lVert\bigg(1-\frac{\mu_{1}(d\zeta)}{\mu_{2}(d\zeta)}\bigg)\Big\rVert^{2}_{\zeta}\mu_{\scriptscriptstyle P}(d\zeta)<\infty\quad\Leftrightarrow\quad\sum_{n\in\mathbb{N}}d(a_{n})\bigg(1-\frac{\mu_{1}(\{a_{n}\})}{\mu_{2}(\{a_{n}\})}\bigg)^{2}<\infty.$
That is, $P_{1}\equiv P_{2}$ on $\sigma_{T}(X)$ if and only if Chow’s (ii) is satisfied.
3 On ML-inference for centered Gaussian processes
Let $T=\{t_{1},t_{2},\dotsc\}$ be countably infinite and introduce $\Theta\subset\mathbb{R}^{p}$. Given $\theta\in\Theta$, let $P_{\theta}$ be a centered Gaussian measure on $\sigma_{T}(X)$ with strictly positive-definite covariance function $R_{\theta}$. We observe that $\sigma_{T}(X)=\sigma\big(\cup_{n=1}^{\infty}\sigma(Y_{n})\big)$, with $\sigma(Y_{n})$ the $\sigma$-filed generated by the random vector $Y_{n}=(X_{t_{1}},\dotsc,X_{t_{n}})$. The set $\Theta$ is regarded as the parameter space and any sequence of random variables $(\hat{\theta}_{n})$ which maximizes the likelihood function over the parameter space is referred to as a sequence of maximum likelihood (ML) estimators. In the present case, the likelihood function is given by $\theta\mapsto p_{\theta}^{n}(Y_{n})$, where $p_{\theta}^{n}$ is defined as in (4). Given $\theta_{0}\in\Theta$, a sequence of ML estimators is said to be strongly consistent for $\theta_{0}$ if
$P_{\theta_{0}}\big(\hat{\theta}_{n}\xrightarrow[]{n\to\infty}\theta_{0}\big)=1.$
It is said to be weakly consistent with limit $\theta_{0}$, if $(\hat{\theta}_{n})$ converges to $\theta_{0}$ in probability $P_{\theta_{0}}$. It turns out that the separability condition
$\theta_{1}\neq\theta_{2}\ \Rightarrow\ P_{\theta_{1}}\perp P_{\theta_{2}}\text{ on $\sigma_{T}(X)$,}\quad\theta_{1},\theta_{2}\in\Theta,$ (20)
relates to the feasibility to estimate $\theta_{0}$ consistently. In particular, a violation of (20) can lead to inconsistent ML estimators [zhang, zhang2005towards, anderes2010consistent, bevilacqua2019estimation, Bachoc2021]. Building on a classical likelihood argument [wald], it is possible to obtain strongly consistent ML covariance parameter estimators for families of orthogonal Gaussian distributions (cf. [zhang, hediger]). The argument relies on the fact that under the separability condition (20), the Radon-Nikodym derivative $\nicefrac{{p_{\theta}^{n}(Y_{n})}}{{p_{\theta_{0}}^{n}(Y_{n})}}$ converges to zero with $P_{\theta_{0}}$ probability one whenever $\theta\neq\theta_{0}$ (cf. Theorem 1 on p. 442 in [StochasticProcesses]). Then, under suitable conditions on the parameter space and the likelihood function, the latter convergence can be strengthened to a $P_{\theta_{0}}$ a.s. uniform convergence on $\Theta$ outside a small neighborhood of $\theta_{0}$. Regarding weak consistency, the reader is referred to [sweeting1980uniform, mardia1984maximum, bachoc2014asymptotic]. In particular, in the latter two, weakly consistent estimators are obtained under the assumption that the distance between coordinates from $T$ is uniformly bounded away from zero.
Acknowledgments.
I thank Serge Cohen — our exchange at the University of Toulouse was a great pleasure.